# Archetype-Aware Predictive Autoscaling with Uncertainty Quantification for Serverless Workloads on Kubernetes

### v1

@@ Abstract
High-performance extreme computing (HPEC) platforms increasingly adopt serverless paradigms, yet face challenges in efficiently managing highly dynamic workloads while maintaining strict performance requirements. We present AAPA (Archetype-Aware Predictive Autoscaler), a novel ML-driven autoscaling system that leverages weak supervision to automatically classify 300,000 workload samples into four distinct archetypes (PERIODIC, SPIKE, RAMP, STATIONARY_NOISY) with 99.8% accuracy, then applies archetype-specific scaling strategies with uncertainty quantification. Evaluation on Azure Functions traces reveals that AAPA achieves up to 50% reduction in SLO violations and 40% improvement in response times compared to Kubernetes HPA, but at 2-8× higher resource consumption—demonstrating that ML-driven approaches shift rather than eliminate the fundamental performance-cost tradeoff in HPEC serverless environments.
@@ I Introduction
Serverless computing has emerged as a transformative paradigm, promising automatic resource management and pay-per-use pricing. Analysis of Azure Functions reveals workload heterogeneity spanning over 8 orders of magnitude in invocation rates [1], presenting significant challenges for resource allocation. Current autoscaling mechanisms in platforms like Kubernetes often employ one-size-fits-all strategies that fail to adapt to diverse workload patterns, leading to either Service Level Objective (SLO) violations or resource over-provisioning.
The core challenge lies in the fundamental mismatch between static scaling policies and dynamic workload characteristics. Social media applications exemplify this heterogeneity: user-generated events create sudden traffic spikes requiring aggressive scaling, while backend analytics exhibit predictable periodic patterns amenable to proactive allocation. A uniform scaling strategy cannot efficiently handle both scenarios, leading to either SLO violations from under-provisioning or cost inefficiency from over-provisioning. Recent industrial deployments combining ML traffic prediction with cloud autoscaling have demonstrated 35% peak reduction through 2-minute ahead forecasting [2], highlighting the potential of workload-aware approaches.
We present AAPA (Archetype-Aware Predictive Autoscaler), a novel autoscaling system that addresses this challenge through three key contributions:
1. Weak supervision framework for automatic workload characterization: A sliding-window feature extraction pipeline that processes serverless traces using 10 programmatic labeling functions to automatically classify 300K time windows from Azure Functions data into four distinct archetypes (spike, periodic, ramp, stationary-noisy) with 99.8% accuracy.
2. Archetype-aware scaling strategies with uncertainty quantification: Machine learning models that predict workload archetypes and apply differentiated scaling policies, incorporating prediction confidence to adopt conservative strategies when uncertainty is high.
3. Comprehensive evaluation framework with open-source implementation: The Resource Efficiency Index (REI) metric that balances SLO satisfaction, resource efficiency, and scaling stability, validated through extensive simulation experiments. We open-source our simulation framework and analysis code at https://github.com/GuilinDev/aapa-simulator.
Evaluation on Azure Functions traces shows AAPA achieves 20-40% SLO violation reduction for spike workloads through warm pools and pre-scaling, while reducing costs 15-25% for periodic workloads. Uncertainty-aware mechanisms prevent mis-scaling, improving stability.
The key insight is that workload diversity should be embraced rather than abstracted away. By explicitly modeling different workload archetypes and their unique scaling requirements, we can achieve better performance-cost trade-offs than generic solutions. This work contributes to the growing field of AI for Systems, demonstrating how machine learning with proper uncertainty quantification [3] can enhance traditional systems problems.
While our evaluation uses Azure Functions traces, the challenges of unpredictable bursts and heterogeneous workloads are equally critical in HPEC domains. Defense systems processing satellite imagery experience sudden spikes when detecting events of interest, scientific simulations generate periodic checkpoints, and real-time sensor networks from IoT deployments create ramp patterns as devices come online. These HPEC workloads demand both high performance and efficient resource utilization—requirements that AAPA’s archetype-aware approach directly addresses. By providing sub-second classification and uncertainty-aware scaling, our system enables HPEC platforms to maintain strict latency SLOs while optimizing resource allocation across diverse computational patterns.
The remainder of this paper is organized as follows: Section II reviews related work, Section III details our methodology, Section IV describes the experimental setup, Section V presents results, and Section VI concludes with future directions.
@@ II-A Serverless Workload Characteristics
Shahrad et al. [1] provided the first comprehensive characterization of serverless workloads through analysis of Azure Functions traces. Their study revealed significant heterogeneity in function execution patterns, with invocation rates varying by over 8 orders of magnitude. This diversity motivates our archetype-based approach.
Table I summarizes common serverless workload patterns identified in recent studies.
@@ II-B Autoscaling in Container Orchestrators
Kubernetes Horizontal Pod Autoscaler (HPA) represents the current state-of-practice for container autoscaling [4]. HPA uses a reactive control loop that scales based on observed metrics (CPU, memory, custom metrics) compared against target thresholds. While simple and robust, this approach suffers from inherent lag in responding to rapid workload changes.
Recent enhancements include predictive scaling capabilities [5], which use time-series forecasting to anticipate future load. However, these systems typically apply uniform prediction models regardless of workload characteristics, missing opportunities for specialized strategies. Recent work in HPC environments has explored adaptive resource management [6, 7], highlighting the importance of workload-aware scheduling in extreme-scale systems.
@@ II-C Machine Learning for Autoscaling
The application of ML to autoscaling has gained significant attention. Table II compares our approach with recent ML-driven autoscaling systems.
@@
Our work differs by explicitly modeling workload diversity through archetypes and incorporating uncertainty quantification, enabling immediate deployment without extensive per-application training. Recent RL-based systems such as AWARE significantly reduce SLO violations by combining horizontal and vertical actions [9], while uncertainty-aware frameworks like MagicScaler dynamically inflate predictions to hedge against forecast error [10]. Our work differs by coupling workload archetyping with calibrated confidence, allowing strategy-level decisions to exploit domain structure. Additionally, serverless-specific autoscalers using DQN on Knative platforms have shown promise for rapid convergence [11], but lack the workload-aware differentiation central to our approach.
@@ II-D Weak Supervision for Systems
Weak supervision has emerged as a powerful paradigm for creating labeled datasets without manual annotation [12]. While primarily used in NLP and computer vision, its application to systems problems remains underexplored. We demonstrate that weak supervision is particularly well-suited for workload characterization, where domain experts can encode heuristics as labeling functions.
@@ III Methodology
III-A System Overview
AAPA comprises three components: (1) feature extraction pipeline processing invocation traces, (2) workload classifier with uncertainty estimates, and (3) adaptive autoscaler applying archetype-specific strategies (Figure 1).
Figure 1: AAPA architecture.
III-B Workload Characterization Pipeline
III-B1 Sliding Window Feature Extraction
We process invocation time series using 60-minute sliding windows (10-minute stride), extracting 38 features across statistical, time-domain, and frequency-domain categories. This approach aligns with findings that different applications require different metrics [13]. Full feature list: https://github.com/GuilinDev/aapa-simulator.
III-B2 Weak Supervision for Workload Labeling
Manual labeling of hundreds of thousands of time windows is infeasible. Instead, we employ programmatic weak supervision [14] with ten domain-specific labeling functions covering spike detection (kurtosis ¿ 10, max-to-median ratio ¿ 20), periodicity (spectral entropy ¡ 0.5, autocorrelation ¿ 0.6), ramp patterns (strong linear trends), and stationary-noisy patterns. We aggregate LF outputs using majority voting, with agreement levels providing natural confidence scores. Complete labeling logic is detailed in our supplementary materials.
III-C Archetype-Aware Predictive Autoscaling
III-C1 Workload Classification
We train a LightGBM classifier on the weakly-supervised dataset to predict workload archetypes from window features. The model outputs both predicted labels and class probabilities, which serve as confidence scores for uncertainty-aware scaling. This systematic incorporation of prediction uncertainty addresses a key limitation in existing autoscalers, enabling adaptive strategy selection based on confidence levels—aggressive scaling when predictions are certain, conservative approaches when uncertain.
III-C2 Differentiated Scaling Strategies
Each archetype employs a tailored scaling strategy as shown in Table III:
@@ III-C2 Differentiated Scaling Strategies
Spike strategy maintains low utilization targets and warm pools to absorb sudden load. Periodic strategy uses time-series forecasting (Holt-Winters) to pre-scale before anticipated peaks. Ramp strategy extrapolates trends to stay ahead of gradual changes. Stationary strategy emphasizes stability with longer cooldowns to prevent oscillation.
III-C3 Uncertainty-Aware Scaling
We calibrate prediction probabilities using beta calibration [15] to obtain reliable confidence scores $c\in[0,1]$. Algorithm 1 shows how uncertainty modulates scaling decisions:
@@ III-C3 Uncertainty-Aware Scaling
0:  Confidence score $c\in[0,1]$, base parameters
0:  Adjusted scaling parameters
@@
2:  $cpu_{adj}\leftarrow cpu_{target}(1-0.2(1-c))$
3:  $cool_{adj}\leftarrow cool_{base}\times m$cooldown}
4:  $rep_{adj}\leftarrow\lceil rep_{base}\times m\rceil$replicas}
@@
This ensures conservative behavior when predictions are uncertain, preventing aggressive mis-scaling.
III-D Resource Efficiency Index (REI)
We propose REI as a unified metric balancing three objectives:
$\text{REI}=\alpha\cdot S_{SLO}+\beta\cdot S_{eff}+\gamma\cdot S_{stab}$
where $S_{SLO}$ is SLO satisfaction rate (1 - violation_rate), $S_{eff}$ is resource efficiency (1 / normalized_pod_minutes), and $S_{stab}$ is scaling stability (1 / scaling_actions). We use default weights $\alpha=0.5$, $\beta=0.3$, $\gamma=0.2$, prioritizing performance while considering cost and stability. Unlike simpler Cost-per-SLO metrics, REI captures the multi-dimensional nature of autoscaling tradeoffs and prevents pathological cases where minimizing one dimension severely impacts others.
IV Experimental Evaluation
@@ IV-A Dataset and Preprocessing
We use the Azure Functions dataset [1] described in Section II, focusing on HTTP-triggered functions that represent user-facing workloads with stringent SLO requirements.
Our preprocessing pipeline:
  1. 1.
    Filter functions with $\geq$1000 total invocations to focus on active workloads
  2. 2.
    Extract 60-minute sliding windows with 10-minute stride
  3. 3.
    Apply weak supervision labeling, yielding  300K labeled windows
  4. 4.
    Split data: days 1-9 for training, days 10-11 for validation, days 12-14 for testing
The resulting dataset exhibits natural class imbalance reflecting real-world distributions: 35% spike, 30% stationary-noisy, 25% periodic, and 10% ramp patterns.
IV-B Implementation Details
AAPA is implemented in Python using LightGBM for classification. The autoscaling simulation uses SimPy to model Kubernetes dynamics with 30-second pod startup times, CPU-based scaling with 1-minute metric aggregation, and FIFO request queuing. All experiments ran on a single NVIDIA RTX 3080 GPU (10GB VRAM) with Ubuntu 24.04 LTS; simulations averaged 7 minutes per workload-day, enabling rapid iteration. Classification latency averaged 2.3 ms per window—negligible at Kubernetes time-scales.

### v3

# AAPA: An Archetype-Aware Predictive Autoscaler with Uncertainty Quantification for Serverless Workloads on Kubernetes

@@ Abstract
Serverless platforms such as Kubernetes are increasingly adopted in high-performance computing, yet autoscaling remains challenging under highly dynamic and heterogeneous workloads. Existing approaches often rely on uniform reactive policies or unconditioned predictive models, ignoring both workload semantics and prediction uncertainty. We present AAPA, an archetype-aware predictive autoscaler that classifies workloads into four behavioral patterns—SPIKE, PERIODIC, RAMP, and STATIONARY—and applies tailored scaling strategies with confidence-based adjustments. To support reproducible evaluation, we release AAPAset, a weakly labeled dataset of 300,000 Azure Functions workload windows spanning diverse patterns. AAPA reduces SLO violations by up to 50% and lowers latency by 40% compared to Kubernetes HPA, albeit at 2–8$\times$ higher resource usage under spike-dominated conditions. To assess trade-offs, we propose the Resource Efficiency Index (REI), a unified metric balancing performance, cost, and scaling smoothness. Our results demonstrate the importance of modeling workload heterogeneity and uncertainty in autoscaling design.
[…]
https://github.com/GuilinDev/aapa-simulator
[…]
@@ I Introduction
Serverless computing has emerged as a key abstraction in both cloud-native and high-performance extreme computing (HPEC) platforms [1], offering on-demand resource scaling, simplified deployment, and cost-efficiency. In practice, serverless applications exhibit a wide spectrum of workload patterns—from sudden traffic bursts triggered by viral user activity, to predictable periodic loads from batch jobs and analytics pipelines, to gradual ramp-up behaviors in IoT deployments [2, 3, 4]. Existing autoscaling mechanisms often fail to accommodate this heterogeneity, resulting in either service-level objective (SLO) violations due to underprovisioning or unnecessary cost inflation from overprovisioning [5]. This raises a fundamental systems challenge: how to allocate just enough resources to meet stringent performance targets under highly dynamic and diverse workloads [6, 7, 3, 8].
The default Kubernetes HPA relies on reactive control loops and static threshold policies that treats all workloads uniformly, leading to sluggish responses and frequent under- or over-provisioning. While recent ML-based approaches [5, 9] improve responsiveness through traffic prediction, they typically train global models that fail to differentiate workload patterns and lack uncertainty modeling—resulting in mis-scalings under ambiguous signals. These limitations underscore the need to incorporate workload semantics and prediction confidence into autoscaling.
A major obstacle to advancing autoscaling research is the lack of labeled datasets for workload pattern recognition. While traces like Azure functions [1] provide abundant data, they lack semantic annotations. We address this gap with AAPAset: 300,000 serverless windows automatically labeled into four archetypes (SPIKE, PERIODIC, RAMP, STATIONARY) using weak supervision on statistical features.
Building on AAPAset, we present AAPA (Archetype-Aware Predictive Autoscaler), a scalable machine learning–based framework that classifies serverless workloads into behavioral archetypes and applies confidence-weighted scaling strategies. Our evaluation shows AAPA reduces SLO violations by up to 50% and tail latency by 40% compared to Kubernetes HPA, while achieving steady gains on our proposed Resource Efficiency Index (REI)-a unified metric balancing performance, cost, and scaling smoothness. While AAPA incurs 2–8$\times$ higher resource usage under high-variance scenarios, this trade-off is justified for latency-sensitive services requiring strict performance guarantees.
Our contributions are as follows:
  • We introduce AAPAset, a labeled dataset of 300K serverless workload windows, weakly supervised into 4 archetypes: SPIKE, PERIODIC, RAMP, STATIONARY.
  • We develop AAPA, an uncertainty-aware autoscaler that dynamically selects archetype-specific scaling strategies based on calibrated confidence scores, reducing SLO violations by up to 50% and latency by 40%.
  • We propose the Resource Efficiency Index (REI), a unified metric that balances performance, cost, and scaling smoothness for comprehensive autoscaler evaluation.
@@ II-A Serverless Workload Characteristics
Shahrad et al. [1] characterized Azure Functions traces, revealing 8-order-of-magnitude variations in invocation rates. Recent studies [10, 11] confirm this heterogeneity across providers, motivating our archetype-based approach. Table I summarizes common serverless workload patterns.
@@ II-B Autoscaling in Container Orchestrators
Kubernetes HPA represents the current state-of-practice for container autoscaling [12]. HPA uses a reactive control loop that scales based on observed metrics (CPU, memory, custom metrics) compared against target thresholds. While simple and robust, this approach suffers from inherent lag in responding to rapid workload changes.
Recent advances in predictive autoscaling leverage time-series forecasting to anticipate future load [13]. However, these approaches typically apply uniform models across all workloads, neglecting behavioral differences that could inform more targeted scaling decisions. In parallel, research in high-performance computing (HPC) systems has emphasized workload-aware resource management [14, 15], reinforcing the broader importance of tailoring policies to workload characteristics in dynamic, large-scale environments.
@@ II-C Machine Learning for Autoscaling
A growing body of work applies machine learning to improve autoscaling decisions. These approaches leverage time-series forecasting, reinforcement learning (RL), or control-theoretic models to enhance responsiveness and cost-efficiency. Table II compares our method with representative ML-based autoscaling frameworks along three dimensions: workload differentiation, predictive uncertainty modeling, and per-application training requirements.
@@
Our approach differs in two fundamental ways. First, it explicitly models workload diversity through archetype classification, enabling generalization without retraining. Second, it incorporates calibrated uncertainty into scaling decisions, allowing the system to adjust strategies based on prediction confidence. Prior methods often treat workloads uniformly—FIRM and AWARE use RL-based controllers without semantic differentiation, MagicScaler [18] inflates predictions to hedge against errors but lacks workload semantics, and DQN-Knative [19] demonstrates fast convergence but ignores workload-specific behavior, which is central to our framework.
@@ II-D Weak Supervision for Systems
Weak supervision enables scalable dataset labeling without manual annotation [20]. Though widely used in NLP and vision, it remains underutilized in systems research. We show it is well-suited for workload characterization, where domain heuristics (e.g., periodicity detection via autocorrelation) can be encoded as labeling functions to annotate large-scale traces. Our method leverages statistical and temporal patterns to generate high-quality labels at scale, illustrating the potential of weak supervision in system-level data contexts.
@@ III Methodology
As shown in Fig. 1, AAPA consists of three integrated components that address the key challenges of heterogeneous workload management: (A) AAPAset, a weakly supervised dataset of workload archetypes from Azure Functions traces; (B) Archetype-Aware Autoscaling, which pairs a calibrated classifier with archetype-specific strategies; and (C) Evaluation and Feedback, which uses a custom REI metric to assess performance and guide refinement.
Figure 1: The overview of AAPA architecture: (A) AAPAset Construction, (B) AAPA, and (C) Evaluation and Feedback.
III-A AAPAset Construction
To support research in archetype-aware autoscaling, we construct AAPAset, a labeled dataset derived from real-world Azure Functions traces [1]. These traces capture the invocation patterns of 5,234 distinct HTTP-triggered functions over a 14-day period, recorded at 1-minute resolution. We retain functions with over 1,000 invocations to exclude ephemeral and cold-start–dominated cases, focusing on HTTP-triggered workloads where autoscaling is most latency-critical.
III-A1 Sliding Window Features
To capture workload dynamics, we segment each function’s time series into overlapping windows using a 60-minute window size and a 10-minute stride, resulting in approximately 300K samples. From each window, we extract 37 hand-engineered features across three categories [^1] Feature definitions and preprocessing code are publicly available: https://github.com/GuilinDev/aapa-simulator.: (1) Statistical features such as mean, standard deviation, skewness, kurtosis, and percentiles; (2) Time-domain features including peak-to-mean ratio, slope, trend direction, and autocorrelation; and (3) Frequency-domain features such as spectral entropy, dominant frequency, and energy in top bands. These features are informed by prior workload modeling research [21] and empirically validated for class separability.
III-A2 Weak Supervision
Manually labeling over 300K windows is infeasible. We instead adopt weak supervision [22], defining ten labeling functions (LFs) that capture domain knowledge to detect temporal patterns. Each LF is a Boolean rule over extracted features:
  • SPIKE: high kurtosis ($>$10) and max-to-median ratio ($>$20)
  • PERIODIC: low spectral entropy ($<$0.5) and high autocorrelation ($>$0.6)
  • RAMP: consistent linear slope (R2 $>$0.8)
  • STATIONARY: low standard deviation and low spectral energy variation
Each window is labeled via majority vote across applicable LFs, with tie cases resolved using fallback heuristics. Agreement strength defines a soft confidence score. This yields a four-class dataset—SPIKE, PERIODIC, RAMP, STATIONARY. Label quality is validated by training a LightGBM classifier, which achieves 99.8% test accuracy, indicating high intra-class consistency.
III-B Archetype-Aware Predictive Autoscaling (AAPA)
AAPA integrates a calibrated workload classifier with archetype-specific control policies for adaptive resource management. It comprises: (1) a confidence-calibrated workload classifier, (2) a set of archetype-specific scaling strategies, and (3) a dynamic adjustment mechanism that adapts scaling behavior based on classification uncertainty.
III-B1 Workload Classification with Uncertainty
We formulate archetype prediction as a supervised classification task, where each input corresponds to a 37-dimensional feature vector extracted from a time window. We adopt LightGBM for its robustness to heterogeneous features, missing values, and data sparsity—common in production serverless traces. The classifier outputs both the predicted label $\hat{y}\in{\textsc{SPIKE},\textsc{PERIODIC},\textsc{RAMP},\textsc{STATIONARY}}$ and the associated probability vector $p=(p_{1},...,p_{4})$.
To ensure the reliability of these confidence scores, we apply beta calibration [23], a post-hoc calibration method that fits class-specific sigmoid functions to the predicted probabilities. This produces a scalar confidence value $c\in[0,1]$, which modulates the aggressiveness of subsequent autoscaling decisions. For example, a confident SPIKE prediction prompts aggressive provisioning, while an uncertain RAMP detection leads to more conservative behavior.
III-B2 Archetype-Specific Scaling Strategies
Each archetype exhibits distinctive temporal patterns, necessitating customized autoscaling strategies. Our archetype-specific policies are:
  • SPIKE: pre-warming, low target CPU (30%), long cooldowns.
  • PERIODIC: predictive scaling (e.g., Holt-Winters), short cooldowns, high CPU targets (75%).
  • RAMP: trend-following extrapolation, medium CPU targets (60%), moderate cooldowns.
  • STATIONARY: conservative scaling with stable resource use and higher CPU thresholds (55%).
Table III summarizes the base parameters for each policy. Notably, the classifier governs both the scaling logic and the tuning of its parameters, in contrast to conventional heuristic-based HPA configurations.
@@ III-B2 Archetype-Specific Scaling Strategies
III-B3 Uncertainty-Aware Scaling Adjustment
To mitigate the risks of over- or under-scaling under uncertain predictions, we introduce a dynamic adjustment mechanism (Algorithm 1) that modifies autoscaling parameters based on the classifier’s confidence score $c\in[0,1]$. When confidence is low, the algorithm increases a margin multiplier $m$ to lengthen cooldown durations, increase replica counts, and modestly reduce the CPU utilization target. This yields more conservative scaling behavior, allowing the system to hedge against misclassifications. All base parameters—$cpu_{target}$, $cool_{base}$, and $rep_{base}$—are defined per archetype in Table III.
@@ III-B3 Uncertainty-Aware Scaling Adjustment
0:  Confidence $c\in[0,1]$, base parameters from Table III
0:  Adjusted parameters $(cpu_{adj},cool_{adj},rep_{adj})$
@@
2:  $cpu_{adj}\leftarrow cpu_{target}\cdot(1-0.2(1-c))$
3:  $cool_{adj}\leftarrow cool_{base}\cdot m$
4:  $rep_{adj}\leftarrow\lceil rep_{base}\cdot m\rceil$
@@
This mechanism ensures that lower-confidence predictions yield more conservative scaling behavior—for example, through longer cooldowns or higher CPU thresholds—thus reducing oscillations and improving system stability under uncertainty.
III-C Evaluation and Performance Feedback
III-C1 Resource Efficiency Index (REI)
To assess and refine autoscaling behavior, AAPA incorporates the REI as both a unified evaluation metric and a feedback signal for runtime adaptation. REI captures the multi-objective nature of autoscaling—balancing performance, efficiency, and stability—within a single interpretable score. Formally:
$\text{REI}=\alpha S_{\text{SLO}}+\beta S_{\text{eff}}+\gamma S_{\text{stab}},\quad\text{where }\alpha+\beta+\gamma=1$ (1)
Each component represents a core operational goal:
[…]
  • $S_{\text{eff}}$: resource efficiency, measured via average CPU utilization and replica-minutes (normalized by workload demand).
  • $S_{\text{stab}}$: system stability, penalizing aggressive oscillations and persistent over-/under-provisioning.
We adopt default weights $\alpha=0.5$, $\beta=0.3$, and $\gamma=0.2$, which emphasize SLO adherence in latency-sensitive deployments. These weights are tunable to reflect alternate priorities such as energy savings or infrastructure stability.
III-C2 Performance Feedback Loop.
Beyond offline evaluation, AAPA employs REI as a feedback signal to support dynamic policy refinement. For instance, sustained REI degradation may trigger:
  • Recalibration of classifier confidence thresholds.
  • Adjustments to archetype-specific scaling parameters (e.g., CPU target or cooldown time).
  • Re-training or fine-tuning of the workload classifier with recent traces.
This feedback loop enables AAPA to adapt to workload drift and infrastructure changes over time, promoting long-term autoscaling robustness. REI thus plays a dual role in AAPA—as both a comparative evaluation tool and a runtime signal for continual improvement.
IV Experimental Setup
We evaluate AAPA across three dimensions: service performance, resource efficiency, and system stability.
@@ IV-A Dataset and Preprocessing
We evaluate on the AAPAset dataset [^1] https://github.com/GuilinDev/aapa-simulator/tree/main/dataset introduced in Section III-A, which comprises 60-minute invocation windows extracted from real-world HTTP-triggered Azure Functions. The dataset is split temporally—days 1–9 for training, 10–11 for validation, and 12–14 for testing—with natural class imbalance in the test set: 35% SPIKE, 30% STATIONARY, 25% PERIODIC, and 10% RAMP.
IV-B Implementation and Simulation
AAPA is implemented in Python using LightGBM for workload classification and SimPy for discrete-event autoscaling simulation. The simulator emulates key Kubernetes behaviors, including pod startup latency (2 seconds), 1-minute metric aggregation, and FIFO request queuing.
Experiments are executed on a workstation with an NVIDIA RTX 3080 GPU (10GB VRAM) running Ubuntu 24.04. Each 1-day workload simulation completes in under 7 minutes, and classification latency averages 2.3ms per window—negligible compared to Kubernetes control intervals.
Each simulation replays a full day of invocation traces. We initialize all workloads with 2 replicas (max = 100), assigning each pod 1000 millicores of CPU and 256MB memory. All experiments are repeated for 5 independent trials per strategy, and results are reported with 95% confidence intervals.